OpenStudy (anonymous):

Solve the following quadratic equation: (x + 3)2 = 16. A. x = -3, -4 B. x = 1 C. x = -1, 7 D. x = 1, -7

OpenStudy (atlas):

There are two ways to go about solving it

OpenStudy (atlas):

I will go with the easy approach first: Which quantity when sqaured will give 16?

OpenStudy (anonymous):

4

OpenStudy (atlas):

4 is right! Any other number you can think which when squared gives 16

OpenStudy (anonymous):

not that i can think of

OpenStudy (atlas):

what abt -4.........what happens when you square (-4)

OpenStudy (anonymous):

-16

OpenStudy (atlas):

nope

OpenStudy (anonymous):

that what my calculator shows

OpenStudy (atlas):

squaring -4 also gives 16 (-4)^2 =16

OpenStudy (atlas):

put brackets around -4 and then see in your calculator

OpenStudy (anonymous):

gotcha -x-=+ i should have remembered that

OpenStudy (atlas):

can you solve the question now

OpenStudy (atlas):

oh and there is another way of doing it (a bit longer) ...........expand the expression (x+3)^2 using the formula (a+b)^2 = a^2 +2ab +b^2 and then solve the equation